FEL Upgrade Project Weekly Brief
February 17-21, 2003
Commissioning highlights this week include: (1) first power tests on the first recirculation arc magnetic string; (2) successful power tests of the wiggler/optical klystron assembly and (3) completion and successful check-out of all internal wiring on the second optical cavity (the outcoupler).

Commissioning Activities:
Beam Transport:
We have a few flanges to make up and one piece of straight vacuum pipe in the return leg remaining in the first light configuration. These will be completed as we are able to catch vacuum technician labor over the next week. Good progress was made in re-attaching the Purcell gap shims on the large recirculation magnet for the second arc (GY)
The wiring was completed and successfully checked out for the second optical cavity (the outcoupler). By late today (Friday), the top hat assembly was being installed and the system will go under vacuum. The other optical cavity (the high reflector) has been pumping all week.
All control software for the optical klystron (wigglers/dispersion section) has been fully exercised and the full assembly has been successfully power tested. One minor problem with a shunt remains on the punch list.
Photoelectron Gun:
The gun ceramic stack assembly was put back under vacuum without the electrode structure and is undergoing a hot nitrogen purge prior to high voltage system tests. The primary electrode set is being repolished. Thermal tests on the filed emission suppression coating on test electrodes passed thermal stress test to 300C.

Special Meeting Notices:
FEL Upgrade Planning Meeting: March 3, 2003
At Jefferson Lab (ONR/DOE invitees only)

FEL Users and Laser Processing Consortium Workshop: March 19-20, 2003
Our annual users meeting will be held on March 19-20 at the CEBAF Center Auditorium at Jefferson Lab. This meeting is open to anyone interested in past and planned future use of the FEL User Facility for basic, applied and mission related R&D. The preliminary agenda for this meeting will be posted on the JLab WEB site next week.

FEL Upgrade Cost and Schedule Review: March 24, 2003
At Jefferson Lab (ONR invitees only)

FEL IR/UV Upgrade Semiannual Review: April 9-10, 2003
At Jefferson Lab (ONR/AFRL/DOE invitees only)

We prepared information for next week’s planning meeting with ONR on cost, schedule and technical issues for 100 kW FEL systems along with updates on our progress on the 10 kW system.

Information was sent to our collaborators and stakeholders in the Applied Research Center consortium because of indications that significant cuts in state funding for this Center were being proposed by the General Assembly in the final week of the 2003 session (which ends tomorrow). The local newspaper (The Daily Press) published a very strong endorsement of the Center on Thursday’s (Feb. 20) editorial page.

The preliminary program for the March 19-20 Laser Processing Consortium/ FEL User’s workshop was sent to our distribution list and posted on JLab’s web site on Tuesday.

We finished our presentation to this weekend’s DOE BES review of new scientific user facilities.

WBS 4 (Injector):
We assembled the gun ceramic stack assembly without the electrode structure and tested the assembly to 50 kV prior to any vacuum processing. A hot nitrogen purge was initiated for vacuum conditioning while the electrode structure is being prepared. The primary electrodes had their field emission coating stripped and they are being repolished according to our standard recipe. They will then be cleaned and re-installed in the gun. In parallel, we are continuing our analysis of why the coating failed, testing the coating procedure, and preparing the back-up set of electrodes for re-coating.

Prior reports noted that we were examining all input variables to the coating process. This week we heated two coated support structures that we had previously prepared last summer when we developed the coating process for the support tube geometry. These tubes were heated to 300C for 3 hrs, held at 300C for 3hrs and cooled. They showed no sign of thermal stress or delamination. We are therefore focusing on the cleaning steps prior to the coating process or the nitrogen plasma cleaning step prior to coating deposition. We have now eliminated field stress or thermal stress as the reason for the coating failure, since the coating did not fail in high field areas and the thermal history for the in –situ bakeout is 250C over a much longer time.

The deposition chamber is undergoing a complete plasma cleaning step this week prior to initiating any more coating tests. We are also checking the integrity of our detergent and solvent cleaning solutions.

WBS 6 (RF):
Zone 4 was tested this week and data taken to check the calibration factors. Several problems developed with the hardware and repaired during the testing. New calibrations factors are to be installed next week.

WBS 8 (Instrumentation):
Beam Loss Monitoring System (BLMs): The development of the new VME cards for the BLM system is in its final stages in Lab 5. (1) The "trip history" feature that allows the user to 'see' in which order the trips occurred is functional as well as (2) the 'beam veto' feature which will momentarily mask the BLMs during the macropulse (for allowable beam modes and correct calculated duty cycles) is functional as well. (3) Another 'new' feature of the BLM system is the status into epics of the active signal strength on all channels. Each channel has a warning level and a trip level status associated with it that will tell the operator how much beam loss each channels 'sees' even while the channel is blanked. A more detailed explanation of the EPICS interface is available this week at: http://laser.jlab.org/wbs8. Each card has proven to work well on its own. The task now is to finalize the high end-screens and integrate the system into the MPS. Updates and cleanup of the BLM HV mainframe software and screens are in progress. Modifications to the FSD screens (CAMAC) were submitted to software. The RF system still relies on the old CAMAC based FSD cards, these inputs now go directly into the new VME based FSD system.

Danfysik Super Trim Power Supplies: 12 of the 16 Apogee PS and PS Interface modules were installed into Zones 5 and 1 respectfully. (see above website) All 3 of the VME Power Supply controllers and control software were also installed. The new screens are located under MEDM/ Magnet/ Magnet Commander/ SF Expert. Installation of the 10 AWG wire leads for the Quads in the 4F and 5F was completed. Klixon termination's for Wiggler were completed and effort has continued on the 1st Arc String.

12 Beam viewers have been assembled using remaining parts. Installation of 3 beam viewers continues at 3F12, 4F00 and 4F00A. Fabrication of the fourth stepper motor driver chassis is in progress. 6 additional Beamviewer cameras have been assembled. Parts are available for 4 more which will be assembled as time permits. 4 new cameras are on order and expected any day.

The design for the optical cavity analog breakout chassis is complete. This will readout the TCs, water temps and flow, and provide for buffered analog inputs from the tunnel directly to EPICS. There are 32 channels of 16 bit, differential input analog dedicated to the cavity. Parts for the new Dump Current Buffer Driver boxes have been received. Assembly and testing will begin next week. Drawing revisions for the Dump Water Control System are complete.

WBS 9 (Transport):
Work Arounds
Substitution of IR Demo 180° Bend (DY) for the Upgrade 180° Bend (GY) in the First Arc and installation of a GW beyond the Optical Klystron as a spectrometer magnet
• Grounding connections, thermal protection connections and water cooling connections were made up.
Optical Chicane Dipoles (GW)
• No progress.
Injector Dipoles (DU/DV)
• Jeff Dale completed re-gluing the Purcell gap plates on the second GU and it is ready for magnetic testing.
Arc 180 Degree Dipoles (GY)
• The Jlab Shop personnel are doing a great job on the removal and refurbishment of the Purcell Gap Shims. A flattening plate piece of tooling was made up to prevent the problem from happening again. The pressing tooling from Wang NMR is due on Saturday morning. Gluing of the Purcell gap on the lower core will happen later in the day on Saturday and the job should be ready for removal of fixturing and clean up by Monday morning. The second core should take less time.
• Contrary to last week’s report, Wang was not able to pot the first GY coil because the silicone tubing used in potting was discovered to be faulty. They installed new tubing and are forecasting the potting for Monday Morning. The second coil is ground wrapped and ready for potting.
Arc Bend & Reverse Bend Dipoles (GQ, GX)
• We started the modifications of the remaining two GQ and two GX magnets to get them up to the level of revision (addition of face and field clamp shims and improvement in field clamp adjustment) so that they are ready for measurement when the hall C quadrupole comes off the stepper stand late next week.
• Wang NMR continues making the replacement GX coil for the #2 GX Dipole with the leaking coil.
Trim Quad (QT)
• The two QT quads with the coils that create the GC vertical corrector are ready for installation in the first arc as soon as Neil Wilson’s folks can be freed up to bring them over to the enclosure.
Sextupole (SF)
• The Sextupole "lights" were internally wired up in the second arc.
Octupole (OT)
• The order is ready to be placed when funding is available.
• Hiroyuki Toyokawa tested out several configurations of spare coils to make up the unique corrector set that we need in the optical chicane. None of the configurations made the acceptable quality of field integral over the wide good field region. Robin Wines is trying out several magnetic models of simple solutions to this problem
Beam Line and Vacuum
• The grounding and additional thermal protection in magnets of the 1st Arc was wired up in anticipation of a commissioning test.
• The 180° Bend chamber for the second arc continues to await the receipt of the diagnostic windows from Ceramaseal. The latest projected delivery date is late February.
• The horizontal corrector in the reverse Bend Dipoles of the first Arc were wired up to terminals on the magnet, awaiting being hooked up to the power supplies.
• We are working on installing the remaining 5 corrector magnet sets around the optical Klystron and on installing the one remaining set at the newly installed Beam Position Monitor just after the extraction chicane.

WBS 11 (Optics):
The HR has been under vacuum since the 19th, and is slowly pumping down. RGA scans show no hydrocarbon contamination. We intend to vent the HR vessel with nitrogen late today or this weekend to install the connecting tube, as the vacuum group has been too busy with work on the main accelerator to do this for us. The wiring issue was repaired last Tuesday and we have been doing the vacuum work. This has taken longer than planned because the vault was locked up early in the day to do work on other subsystems. We believe we will make up the slip in schedule. The water chiller/controller for the optical cavity arrived this week and will be moved downstairs the week after next. All mounting hardware for the insertable mirror after the outcoupler arrived, and it will be installed early next week. We are assembling the OCMMS optical hardware.

Other activities:
Two ultrafast oscillators were installed in User Lab 6. We assisted with technical support, and diagnostics.

With the eminent loss of the Mechanical Engineering Group (MEG) support, drawings for the region after the outcoupler were completed and signed. Regretfully, while these engineers and designers are working on other Divisional tasks, they will be unavailable to work on remaining Optics tasks, such as completion of the optical transport or UV FEL. This will introduce a schedule slip of some weeks to months. The Optics Group would like to express our appreciation for the hard work and ingenuity the MEG supplied to complete our tasks. We look forward to the time when we can work together again.